OpenStudy (anonymous):

Find the inverse of the function f(x)=x^{2}-8x+4, where <4. State its domain.

OpenStudy (anonymous):

\[where x \le 4\]

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

can you complete the square ? know how to ?

hartnn (hartnn):

\(f(x)=x^2-8x+4= (x-...)^2-...\)

OpenStudy (anonymous):

so the answer is inversef(x)=4+sqrt{x+12} . domain is x bigger or equal -12 ?

hartnn (hartnn):

to be more precise, \(f^{-1}(x)=4\pm \sqrt{x+12} \) the domain \(x \ge -12 \) is correct.

hartnn (hartnn):

wait, its given that x<4

hartnn (hartnn):

so, it'll be only \(f^{-1}(x)=4 - \sqrt{x+12}\) with same domain.

OpenStudy (anonymous):

how can we understand that it is minus when x<4 ?

hartnn (hartnn):

4+ anything means >4 4- anything means <4

hartnn (hartnn):

does that clear it up ? or you have different doubt altogether ? :P

OpenStudy (anonymous):

a little bit :D because it tells about x but we take the function <4 :/

hartnn (hartnn):

oh, .... domain of f(x) means range of its inverse function \(f^{-1}(x)\) here, x<4 is domain of f(x), so range of \(f^{-1}(x)\) is x<4 , means \(f^{-1}(x)\) has least value of 4, so we discard 4+sqrt {x+12} because that will be >4

hartnn (hartnn):

***range of \(f^{-1}(x)\) is \(f^{-1}(x)\)<4

hartnn (hartnn):

***maximum value of 4

hartnn (hartnn):

sorry for the typos.

OpenStudy (anonymous):

oh. okey. thank youu :)

hartnn (hartnn):

welcome ^_^